### v1

# Human-free Prompted Based Anomaly Detection: prompt optimization with Meta-guiding prompt scheme

Abstract
Pre-trained vision-language models (VLMs) are highly adaptable to various downstream tasks through few-shot learning, making prompt-based anomaly detection a promising approach. Traditional methods depend on human-crafted prompts that require prior knowledge of specific anomaly types. Our goal is to develop a human-free prompt-based anomaly detection framework that optimally learns prompts through data-driven methods, eliminating the need for human intervention. The primary challenge in this approach is the lack of anomalous samples during the training phase. Additionally, the Vision Transformer (ViT)-based image encoder in VLMs is not ideal for pixel-wise anomaly segmentation due to a locality feature mismatch between the original image and the output feature map. To tackle the first challenge, we have developed the Object-Attention Anomaly Generation Module (OAGM) to synthesize anomaly samples for training. Furthermore, our Meta-Guiding Prompt-Tuning Scheme (MPTS) iteratively adjusts the gradient-based optimization direction of learnable prompts to avoid overfitting to the synthesized anomalies. For the second challenge, we propose Locality-Aware Attention, which ensures that each local patch feature attends only to nearby patch features, preserving the locality features corresponding to their original locations. This framework allows for the optimal prompt embeddings by searching in the continuous latent space via backpropagation, free from human semantic constraints. Additionally, the modified locality-aware attention improves the precision of pixel-wise anomaly segmentation.
1 Introduction
In the field of anomaly detection, acquiring annotations for anomalous samples is often challenging. Consequently, prior research has focused on unsupervised learning, which involves training models on large volumes of anomaly-free data to identify deviations from normal features as anomalies. Recently, some approaches have addressed more extreme scenarios where only a few normal samples are available. These methods leverage vision-language pretrained models, such as CLIP, to facilitate few-shot anomaly detection.
Figure 1: The upper part shows previous methods requiring human-designed prompts. The lower part illustrates our automated, data-driven approach using learnable prompts without prior knowledge.
Vision-language models are renowned for their adaptability to various downstream tasks such as classification and segmentation. The remarkable transferability of vision-language models across different computer vision tasks can be attributed to their large-scale pre-training on text-image pairs [3]. This training enables the model to capture relevant semantic features based on the provided prompts, thereby allowing the model to leverage the textual modality to solve visual tasks such as anomaly detection. [8] pioneered the application of VLMs for zero-shot or few-shot anomaly detection. This approach leverages lexical terms related to ’anomaly’ and ’normal’ to identify anomalous visual features in the test samples, providing a foundation for subsequent studies to further advance few-shot anomaly detection methodologies. Prompt engineering is a crucial technique that preserves the robustness and generalization capabilities of CLIP while efficiently adapting it to meet the demands of specific applications. The primary objective of prompt tuning is to obtain an optimal textual embedding that aligns with the target visual image embedding, ensuring that the model accurately captures the nuances of the task at hand.
Previous work [8] proposed a prompt ensemble technique that composes different prompt templates (e.g., ”a photo of a [c] for visual inspection”) and normal-abnormal-relevant prompts (e.g., ”flawless” for normality, ”damaged” for anomaly) to obtain text embeddings aligned with anomaly detection semantics. Recent studies [9, 16] have leveraged prefix prompt tuning (context optimization [15]) to replace static prompt templates with learnable prompts, better catering to anomaly detection tasks. Although prior work has demonstrated the efficacy of prefix prompting, it still necessitates a customized, manually-designed suffix to describe the anomaly type for each class item. This process heavily relies on detailed and specific prior knowledge of the dataset and the types of anomalies that may appear in each product category, as shown in the upper part of Fig 1. This remains a non-trivial task requiring iterative adjustments, rather than automatically finding the optimal textual embedding.
We are motivated by [15], which provides insights that the optimal prompts derived through prompt tuning do not necessarily align with human semantics or intuition. Previous work has often focused on combining various human semantic terms, such as ’anomaly’ and ’defect,’ to generate better embeddings. However, we propose a novel perspective: instead of adhering to human-centric semantics, we should employ data-driven optimization methods, such as gradient-based learning, to derive the most effective textual embeddings. By utilizing these advanced deep learning techniques, we can create more robust and precise embeddings that better capture the nuances needed for anomaly detection and adapt to various datasets without the need for prior knowledge as show in lower part in Fig 1.
However, to optimize the Learnable Normal Prompt (LNP) and Learnable Anomaly Prompt (LAP), samples from both normal and anomalous data are required. In the case of few-shot anomaly setting, only normal samples are available during training. To enable the acquirement of optimal textual embedding for anomaly detection, we need to self-generate anomaly sample to obtain the optimal prompt for better textual embedding. One intuitive method to synthesize anomalous samples is by adding Gaussian noise to the image pixels to mimic anomalies. However, Gaussian noise alone might not guarantee performance in real-world scenarios because such synthetic anomalies often lack the complex and context-specific characteristics of true anomalies.
Additionally, Vision-Language Models (VLMs) are not designed for pixel-wise tasks like segmentation, making them unsuitable for pixel-level anomaly detection. During VLM pre-training, only global features (e.g., class tokens in Vision Transformers) are considered, neglecting localized features that capture fine-grained image details. Consequently, the relationship between text prompts and specific image regions is not optimized, leading to unpredictable results where the target prompt may not align with the target object pixels [12, 10].
In this paper, we propose a human-free prompting anomaly detection framework, ”Meta-prompting Semantic Learning for Anomaly Detection with Locality-aware Attention Image Encoder”. The proposed framework follows the intuition of data-driven methods that leverage back-propagation to actively find the prompt that can generate the optimal textual embedding suited for anomaly detection, instead of passively combining all possible human semantic prompts to meet the requirements of prompt-based anomaly detection. To address the limitation of the absence of anomaly samples for context optimization, we develop the Object-Attention Anomaly Generation Module to synthesize anomaly samples for training. Meanwhile, the proposed Meta-guiding Prompt-tuning Scheme (MPTS) is developed to prevent the learnable prompt from overfitting on synthesized anomalies through iterative gradient-based calibration.
In Meta-Prompt Tuning Scheme (MPTS), we introduce a dynamic meta-prompt that serves as an anchor to balance specificity and generality in anomaly detection via gradient-based calibration. This meta-prompt ensures that learnable prompts remain aligned with meaningful, generalizable concepts while leveraging useful information from synthesized anomalies without deviating from the core goal of anomaly detection. Initially, a general prompt template provided by a large language model (LLM) is embedded with real-world anomaly semantics, serving as the initial meta-prompt. This helps ensure that the learnable prompts do not deviate significantly from true anomalies during fine-tuning. Throughout each training round, the optimized learnable prompt from the previous round serves as the new meta-guiding prompt, iteratively refining the process. Meta-guiding prompts ensure that the learnable prompts resemble real anomalies, enhancing detection effectiveness. Additionally, introducing Gaussian noise diversifies the patterns, improving generalization across various anomaly types. This combination allows the learnable prompts to generalize well to different anomalies while maintaining alignment with real-world characteristics.
To further improve the quality of synthesized anomalous samples, we introduce the ”Object-Attention Anomaly Generation Module”. This module leverages the general knowledge of CLIP to identify the target object in the image and applies noise specifically to that object. By doing so, the synthesized anomalies better reflect practical scenarios, as real-world anomalies usually occur on target objects rather than in the background, which is typically not the focus of anomaly detection.
Additionally, we introduce Locality-aware Attention, which can be adopted by any transformer-based model to mitigate the misalignment of input token features and corresponding output token features. This mechanism restricts attention to local neighborhoods, preserving essential local details and enhancing the alignment between the input features and the output tokens. By focusing on neighboring regions, the locality-aware attention ensures that the feature extraction process maintains the spatial integrity of the image. This addresses the limitations of the vanilla attention mechanism, where distant tokens tend to dilute crucial local information, resulting in the misalignment of locality features between input and output. aThe contributions are listed as follows:
  • We propose a novel paradigm for a prompt-based anomaly detection framework, namely Meta-prompting Semantic Learning for Anomaly Detection with Locality Feature Attention Image Encoder, which optimizes the pre-trained VLM for anomaly detection tasks in a human-free prompting manner.
  • We introduce the Meta-guiding Prompt-tuning Scheme along with the Object-Attention Anomaly Generation Module to enable the optimization of learnable prompts in the absence of anomaly samples. The optimized learnable prompts with the Meta-guiding Prompt-tuning Scheme even exceed manually designed prompts by a significant margin.
  • We propose the Locality-Aware Transformer, which aims to extract a feature map where all locality features align with their original image positions, ensuring pixel-wise anomaly segmentation in the CLIP-based framework.
Figure 2: The overall architecture of Meta-prompting Semantic Learning for Anomaly Detection with Locality Feature Attention Image Encoder
2 Related Work
2.1 Vision-language model
Vision-language models (VLMs) are advanced AI systems that integrate computer vision (CV) and natural language processing (NLP) to understand the commonality of semantics between text embeddings and visual features. These multimodal models are trained on extensive datasets of paired images and text, utilizing techniques like contrastive learning [3] and masked language modeling to map visual and textual data into a shared semantic space. A notable open-source VLM is OPENCLIP [7], which excels in generalization, enabling rapid adaptation to various downstream tasks such as image classification and object detection without task-specific training. Furthermore, [15] proposed context optimization, which can tune learnable prompts with few-shot samples, achieving optimal performance surpassing human-designed prompts with.
2.2 Anomaly detection
Traditional paradigms of anomaly detection [13, 6, 11, 2] typically utilize only normal samples for training anomaly detection models. This approach, known as unsupervised anomaly detection, does not require any anomaly samples during training. In recent years, the research focus has shifted from unsupervised anomaly detection to the more challenging few-shot scenario, where only a limited number of normal samples are available during the training phase. PatchCore [4] employs a memory-efficient technique to capture normal patterns in few-shot data for anomaly inference. RegAD [14] introduces a region-based method that maximizes the information extracted from each example, effectively learning informative features for anomaly inference using only a few normal samples.
Recent studies [8, 9] have further addressed the anomaly detection challenge with multimodal solutions, leveraging the general knowledge embedded in pre-trained vision-language models (VLMs) to perform zero-shot and few-shot anomaly detection with the guidance of textual modality. However, these approaches still rely on human-designed prompts, with their performance dependent on the prompt engineer’s prior knowledge of the specific application scenario. This dependency highlights the need for more advanced, data-driven methods to optimize prompt selection and enhance anomaly detection performance.
3 Proposed Framework
3.1 Problem Definition:
In this paper, we address the few-shot anomaly detection task, where only a few normal samples are available during the training stage. We suppose $D_{\text{train}}=\{(\mathbf{x}_{i},y_{i}=0)\}_{i=1}^{k}$ is the $k$-shot sample given in the training stage ($y_{i}\in\{0,1\}$). Each $\mathbf{x}_{i}\in\mathbb{R}^{H\times W}$ is a normal sample when $y_{i}=0$, and abnormal when $y_{i}=1$.
In our proposed human-free prompt-based anomaly detection framework, we substitute traditional human prompts (”a photo of a {} without defect” as the normality prompt and ”a photo of a defective {} as the abnormality prompt) with learnable prompts. The learnable normality prompt (LNP) and the learnable abnormality prompt (LAP) are composed of a learnable general prefix (LGP) $=\left[G_{1},\ldots,G_{ng}\right]$ and a learnable normal or abnormal suffix (LNS $=\left[N_{1},\ldots,N_{nr}\right]$ or LAS $=\left[A_{1},\ldots,A_{na}\right]$, respectively:
$\begin{cases}\text{LNP}=\left[G_{1},\ldots,G_{ng}\right]\left[\text{cls}\right]\left[N_{1},\ldots,N_{nr}\right]\ \\ \text{LAP}=\left[G_{1},\ldots,G_{ng}\right]\left[\text{cls}\right]\left[A_{1},\ldots,A_{na}\right]\end{cases}$ (1)
The LNP and LAP are leveraged to calculate the anomaly score map $S\in\mathbb{R}^{H\times W\times 1}$ on the feature map $F\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C}$ extracted from the testing sample.
In the inference stage, we have $D_{\text{test}}=\{(\mathbf{x}_{i},y_{i},M_{i})\}_{i=1}^{N}$. The mask $M\in\{0,1\}^{H\times W}$ is the ground truth map for anomaly pixels, where each pixel $m_{j}\in M$ indicates whether the corresponding pixel $x_{j}\in x$ is anomalous ($m_{j}=1$) or normal ($m_{j}=0$).
Our objective is to leverage $D_{\text{train}}$ to tune LAP and LNP such that they can yield an anomaly score map $S$ that aligns with the ground truth anomaly map $M$. This can be represented as:
$\arg\max\text{Sim}(S,M)$ (2)
where $\text{Sim}()$ is a similarity metric that calculates the alignment degree between $S$ and $M$.
3.2 Overview
The overall framework of the proposed Meta-prompting Semantic Learning for Anomaly Detection with Locality Feature Attention Image Encoder is illustrated in Fig 2. The proposed Meta-guiding Prompt-tuning Scheme (Sec 3.3) sets a new paradigm for prompt-based anomaly detection with no human semantic lexicon involved in inference. To address the absence of anomalous samples for tuning prompts, the proposed Object-Attention Anomaly Generation Module (OAGM) (Sec 3.4) produces synthesized anomalous samples for tuning while remaining practical by only applying to items instead of the background. Additionally, to enhance the locality feature extraction, the visual encoder adopts a novel proposed Locality-Aware Transformer (Sec 3.5), which enhances locality feature extraction with neighbor-cls-only attention. The prompts LG and normal/abnormal suffix ( LNS / LAS ) are designed as fully learnable parameters.
3.3 Meta-guiding Prompt-tuning Scheme
Figure 3: The illustration of Meta-guiding Prompt-tuning Scheme
The Meta-guiding Prompt-tuning Scheme is illustrated in Fig 3, which an self-optimized meta-prompt will iterativly update it self and guiding the learning direction of learnable prompt with gradient correction.
To tune LAP and LNP to align with the anomaly detection task, the OAGM (Sec 3.4) will synthesize anomalous samples ${(\mathbf{x^{\prime}}_{i},y_{i}=1,M^{\prime}a)}{i=1}^{k}$, where $M^{\prime}_{a}\in{0,1}^{H\times W}$ is the abnormality binary mask indicating whether each pixel $m_{j}\in M^{\prime}_{a}$ is contaminated with synthesized anomaly ($m_{j}=1$) or not ($m_{j}=0$). A normality binary mask $M^{\prime}_{n}$ can be obtained by:
$M^{\prime}_{n}=1-M^{\prime}_{a}$ (3)
The LNP with the extracted feature map of $\mathbf{x^{\prime}}_{i}$ is dot-multiplied to calculate the normality score map $S_{n}$, indicating whether each pixel in $\mathbf{x^{\prime}}_{i}$ aligns with normality. Similarly, LAP is used to obtain an abnormality score map $S_{a}$. To aggregate the information from both LNP and LAP and produce a probability-like output (bounded from 0 to 1), we concatenate $S_{n}$ and $S_{a}$ and apply Softmax along each element, finally obtaining normalized score maps $S^{\prime}_{n}$ and $S^{\prime}_{a}$. The training objective is to tune LNP such that its embedding perfectly aligns with the normality embedding in $\mathbf{x^{\prime}}$. This involves aligning $S^{\prime}_{n}$ with the normality binary mask $M^{\prime}_{n}$ using BCEloss:
$\mathcal{L}_{\text{ano}}^{\text{LNP}}=-\sum_{j\mid m_{j}=1}\Big{[}$ $M^{\prime}_{n}(j)\log S^{\prime}_{n}(j)$
$+(1-M^{\prime}_{n}(j))\log(1-S^{\prime}_{n}(j))\Big{]}$ (4)
For LAP, the objective is to align LAP with the abnormality embedding in $\mathbf{x^{\prime}}$:
$\mathcal{L}_{\text{ano}}^{\text{LAP}}=-\sum_{j\mid m_{j}=1}\Big{[}$ $M^{\prime}_{a}(j)\log S^{\prime}_{a}(j)$
$+(1-M^{\prime}_{a}(j))\log(1-S^{\prime}_{a}(j))\Big{]}$ (5)
Meta-guiding: Synthesized anomaly samples, while useful for training, may contain artifacts or specific characteristics that are not representative of real-world anomalies. To address this limitation, we propose a meta-guiding scheme to regularize the tuning process with general, domain-agnostic knowledge embedded in pre-trained VLMs.
Initially, we provide general manual prompts as Meta Normality Prompt (MNP) and Meta Abnormality Prompt (MAP) (see appendix), which act as anchors to prevent LAP and LNP from overfitting on artifact anomalies and maintain generalization with a broader concept of anomaly. By calibrating the LNP and LAP gradients from Eq 3.3 and Eq 3.3, respectively, with the divergence loss between MAP/MNP and LAP/LNP, we can ensure that LAP and LNP are optimized toward a task-specific prompt that human semantics cannot achieve while maintaining generalization.
This divergence loss is designed to estimate the output score maps from LAP/LNP ($S^{\prime}_{a}$ and $S^{\prime}_{n}$, respectively) and MAP/MNP (denoted as $S^{M}_{a}$ and $S^{M}_{n}$, respectively). Since the values of the output score maps are already transformed to probability form, we use KL divergence as the distance metric to estimate the divergence. The divergence loss can be written as:
$\mathcal{L}_{\text{div}}^{A}=\sum_{i=1}^{H\times W}\text{KL}(S^{M}_{a}(i)\parallel S^{\prime}_{a}(i))$ (6)
$\mathcal{L}_{\text{div}}^{M}=\sum_{i=1}^{H\times W}\text{KL}(S^{M}_{n}(i)\parallel S^{\prime}_{n}(i))$ (7)
To strike a balance between task-specific objectives and generalization, the meta-guiding scheme will only calibrate the gradient when the gradient $G_{\text{ano}}^{\text{LAP/LNP}}$ from fine-tuning the anomaly loss:
$G_{\text{ano}}^{\text{LAP/LNP}}=\frac{\partial\mathcal{L}^{\text{LAP/LNP}}_{\text{ano}}}{\partial\theta_{\text{LAP/LNP}}}$ (8)
contradicts the gradient from divergence loss $G_{\text{div}}^{\text{LAP/LNP}}$.
$G_{\text{div}}^{\text{LAP/LNP}}=\frac{\partial\mathcal{L}_{\text{div}}^{\text{A/M}}}{\partial\theta_{\text{LAP/LNP}}}$ (9)
The LAP and LNP are optimized with the gradient $G_{\text{ano}}^{\text{LAP/LNP}}$ after being calibrated with Eq 10:
$\begin{cases}G_{\text{ano}}-\lambda\cdot\text{Cos}\left(G_{\text{ano}},G_{\text{div}}\right),&\text{if }\text{Cos}\left(G_{\text{ano}},G_{\text{div}}\right)<0\\ G_{\text{ano}},&\text{otherwise}\end{cases}$ (10)
The meta-guiding scheme ensures that the optimization direction will explore the optimal prompt that aligns better with the anomaly detection task while preventing overfitting on synthesized anomalies.
Based on the intuition that we have obtained a more optimal $\text{LAP}^{t+1}$/$\text{LNP}^{t+1}$ through gradient descent instead of relying on human semantics, we substitute the MAP/MNP with $\text{LAP}^{t+1}$/$\text{LNP}^{t+1}$ as the new Meta prompt for the next meta-epoch training. We iteratively refine the MAP/MNP and LAP/LNP until the loss converges. This iterative substitution strategy balances the need for task-specific adaptation and generalization. By continually referring back to the meta-guiding prompts, the model maintains a connection to the initial general knowledge while fine-tuning the prompts to better detect anomalies.
3.4 Object-attention Anomaly Generation Module
To tune LAP and LNP to align with the anomaly detection task, we propose the Object-Attention Anomaly Generation Module (OAGM). Traditional methods that add Gaussian noise randomly across the entire image are suboptimal because they affect irrelevant background regions, diluting the model’s ability to focus on the target object. The OAGM addresses this limitation by selectively adding noise only to the target object areas, enhancing the model’s ability to detect anomalies.
First, we utilize a pre-trained Vision-Language Model (VLM) to produce the binary mask $\mathbf{M}_{\text{obj}}$ for the target object:
$\mathbf{M}_{\text{obj}}(i,j)=\begin{cases}1&\text{if pixel }(i,j)\text{ belongs to the target object}\\ 0&\text{otherwise}\end{cases}$
Next, Gaussian noise $\mathcal{N}(0,\sigma^{2})$ is selectively added to the pixels identified by $\mathbf{M}_{\text{obj}}$:
$\mathbf{x}^{\prime}(i,j)=\begin{cases}\mathbf{x}(i,j)+\mathcal{N}(0,\sigma^{2})&\text{if }\mathbf{M}_{\text{obj}}(i,j)=1\\ \mathbf{x}(i,j)&\text{if }\mathbf{M}_{\text{obj}}(i,j)=0\end{cases}$
3.5 Locality-aware attention
The vanilla ViT is not suitable for pixel-level tasks due to the feature misalignment issue. We obtain insight from [5] that the feature misalignment issue results from the nature of the original attention mechanism, where each patch (token) feature attends to distant patch features. This leads to local features being fused with irrelevant background features, causing misalignment between input and output features.
Based on this observation, we substitute the vanilla attention with the proposed Locality-Aware Attention (LAA) that restricts each token to attend only to its neighboring token features during feature extraction. To avoid model collapse caused by directly modifying the self-attention mechanism, which would lead to changes in subsequent layers and be amplified through the layers, we adopt the dual-path design from to maintain stable input features by preserving the original path alongside the new path.
The LAA applies a $k$-neighbor exclusive mask $\mathbf{M}_{\text{LAA}}\in\mathbb{R}^{H^{\prime}\times W^{\prime}}$ to mask the attention map (the dot product of $\mathbf{Q}$ and $\mathbf{K}$) in the attention mechanism:
$\text{Attention}_{\text{LAA}}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}+\mathbf{M}_{\text{LAA}}}{\sqrt{d_{k}}}\right)\mathbf{V}$ (11)
where $\mathbf{M}_{\text{LAA}}$ is constructed with the following formula to enable each token to attend only to its $k$-neighbors:
$M_{(i,j),(m,n)}=\begin{cases}0&\text{if }\sqrt{(i-m)^{2}+(j-n)^{2}}\leq k\\ -\infty&\text{otherwise}\end{cases}$ (12)
Each patch can be indexed in the 2D grid as $(i,j)$ where $0\leq i<H^{\prime}$ and $0\leq j<W^{\prime}$. $(i,j)$ and $(m,n)$ are the coordinates of the patches in the original feature map before patchification, and $k$ is the predefined neighbor distance.
4 Experiments
We conduct experiment to validate that Meta-guiding Prompt-tuning Scheme with Object-attention Generation Module can optimize a prompt embedding that outperform human-semantic prompt under scenario that only 1 sample (1-shot) is provided in pixel-level anomaly segmentation. Additionally, we complete ablation experiment to verify the effectiveness of each proposed component.
Dataset In this paper, we conduct experiment on MVTec[1] dataset, which is a benchmark dataset for industrial anomaly detection that provide pixel-level ground truth map indicating the anomaly location in the image. MVtec comprise 15 Categories of item, and each category contain 5-6 sub-categories of anomlay type.
4.1 Comparison to human-semantic prompt
Table 1: Comparison to Human semantic prompt
Class carpet grid leather tile wood bottle cable capsule hazelnut metal_nut pill screw toothbrush transistor zipper mean
Human-semantic prompt 93.98 35.86 90.89 43.21 57.74 82.86 42.54 72.05 84.09 55.98 51.57 70.77 75.43 62.23 50.90 64.67
Ours 99.40 98.45 99.38 96.12 96.18 89.91 87.75 94.51 97.70 75.08 95.92 95.48 97.23 72.23 90.63 92.40
As shown in Table1, we can see that our proposed framework outperform manual design prompt by 28%, which further validate our assumption that optimal prompt for downstream task is not necessary align with human semantic. The human-semantic prompt template and suffix is followed by the literature . We have further conduct ablation study in Sec. 4.2 to verify that the meta-prompt guiding scheme is crucial for learning the optimal prompt embedding.
4.2 Ablation Study
Ablation on Meta-prompt guiding scheme
Table 2: The ablation study of guiding level $\lambda$ for Meta-guiding Prompt-tuning Scheme
Ablation study on MPTS
Class $\lambda=0$ $\lambda=0.5$ $\lambda=1$
carpet 76.27 99.32 99.40
grid 60.51 97.97 98.45
leather 76.84 99.15 99.38
tile 80.87 94.86 96.12
wood 50.06 95.42 96.18
bottle 49.01 84.26 89.91
cable 48.90 81.80 87.75
capsule 87.55 91.84 94.51
hazelnut 77.84 87.74 97.70
metal_nut 52.74 76.77 75.08
pill 67.36 87.33 95.92
screw 78.54 95.27 95.48
toothbrush 52.90 92.59 97.23
transistor 60.55 65.23 72.23
zipper 88.22 93.90 90.63
mean 67.21 89.56 92.40
The ablation study results in Table2 highlight the effectiveness of the Meta-guiding Prompt-tuning Scheme, particularly in enhancing anomaly detection performance. When $\lambda=0$, the model lacks meta-guiding and achieves a mean score of 67.21, indicating limited effectiveness due to overfitting on synthesized anomalies. Introducing meta-guiding with $\lambda=0.5$ significantly improves the mean score to 89.56, demonstrating enhanced generalization and alignment with broader anomaly concepts. The highest performance is achieved with $\lambda=1$, yielding a mean score of 92.40 and near-perfect results in several classes, confirming that the scheme effectively balances task-specific tuning and general knowledge integration. These findings validate the module’s contribution to improving anomaly detection accuracy and robustness.
Ablation on Object-attention Anomaly Generated Module(OAGM)
Table 3: Comparison of Anomaly Detection Performance with and without OAGM
Ablation study on OAGM
Class Without OAGM With OAGM ($\lambda=1$)
carpet 98.36 99.40
grid 96.53 98.45
leather 97.93 99.38
tile 91.14 96.12
wood 89.19 96.18
bottle 64.18 89.91
cable 49.31 87.75
capsule 85.53 94.51
hazelnut 65.62 97.70
metal_nut 72.12 75.08
pill 78.42 95.92
screw 87.06 95.48
toothbrush 77.05 97.23
transistor 66.85 72.23
zipper 75.61 90.63
mean 78.54 92.40
The ablation study in Table3 demonstrates the effectiveness of the Object-attention Anomaly Generation Module (OAGM) across different categories. For texture items like carpet, wood, and tile, the entire image serves as the target, leading to high baseline performance even without OAGM (98.36, 89.19, and 91.14, respectively). OAGM provides slight improvements (99.40, 96.18, and 96.12) due to the inherent nature of these items. Conversely, object-like items such as bottle, cable, and capsule show significant performance boosts with OAGM, increasing from 64.18, 49.31, and 85.53 to 89.91, 87.75, and 94.51, respectively. The overall mean score rises from 78.54 to 92.40, highlighting OAGM’s role in enhancing anomaly detection by focusing on target objects and reducing irrelevant background noise, especially in object-centric categories.
Ablation on Locality-aware Attention
To evaluate the impact of the proposed Locality-Aware Attention (LAA), we conducted an ablation study comparing different configurations of attention mechanisms. Specifically, we compared the original VV-att (V-V attention as described in [10]), QK-LAA (original QK attention with LAA), and VV-LAA (V-V attention combined with LAA). The results are summarized in Table4.
Table 4: Ablation study on Locality-Aware Attention
Ablation study on Locality-Aware Attention
Class VV-att QK-LAA VV-LAA
carpet 98.29 99.29 99.40
grid 95.93 96.93 98.45
leather 98.10 99.10 99.38
tile 94.63 95.63 96.12
wood 94.07 95.07 96.18
bottle 79.83 80.83 89.91
cable 84.58 85.58 87.75
capsule 89.09 90.09 94.51
hazelnut 94.53 95.53 97.70
metal nut 72.64 73.64 75.08
pill 85.12 86.12 95.92
screw 91.77 92.77 95.48
toothbrush 95.55 96.55 97.23
transistor 70.19 71.19 72.23
zipper 87.45 88.45 90.63
mean 89.52 90.52 92.40
From the results, it is evident that LAA enhances performance across all categories. The mean performance of QK-LAA (90.52) surpasses that of VV-att (89.52), demonstrating that the locality constraint improves feature alignment even when applied to the traditional QK attention mechanism. Furthermore, the combination of V-V attention with LAA (VV-LAA) achieves the highest mean performance (92.40), indicating that our LAA not only addresses the feature misalignment issue but also synergizes effectively with advanced attention mechanisms to boost overall performance. This significant improvement underscores the effectiveness of LAA in enhancing pixel-wise anomaly detection tasks.
5 Conclusion
In this paper, we introduced a novel framework, Meta-prompting Semantic Learning for Anomaly Detection with Locality Feature Attention Image Encoder, which optimizes vision-language models for anomaly detection in a human-free manner. Our Meta-guiding Prompt-tuning Scheme and Object-Attention Anomaly Generation Module address the absence of anomaly samples by synthesizing realistic anomalies and iteratively refining prompts via gradient-based calibration. Additionally, the Locality-Aware Transformer preserves local details and ensures alignment between input features and output tokens, enhancing pixel-wise anomaly segmentation. Our extensive experiments demonstrated significant performance improvements over manually designed prompts, showcasing the efficacy of our approach and advancing the field of few-shot and zero-shot anomaly detection.This work opens new avenues for future research in developing more sophisticated anomaly detection models that can adapt to various real-world scenarios without extensive human intervention.